HOME ESR EXPERIMENTS LABS etc
E121 E127 E132 E125 E143 laser_cooling_2021 E142 Ê128 E146 E0052 E0018 E0028 E0038 E0075 E0203_E0018
  masses & isomeres 2025  Not logged in ELOG logo
Message ID: 63     Entry time: Wed Apr 2 04:59:50 2025
Author: Zac, Chris, Carlo 
Category: General 
Subject: Receiving injections but not seeing anything on RSAs 
Starting at around 04:09 we don't see anything on RSAs. RSAs are still triggering and writing files, but only see noise. Lassie indicates beam current in the ESR

We called HKR, they said they will investigate, awaiting response.


last file that contains data: RSA01-2025.04.02.04.09.27.044.tiq_spectrogram.npz
file from which only noise is contained: RSA01-2025.04.02.04.09.31.905.tiq_spectrogram.npz 

At 04:09: drop of beam in the SIS as shown in screenshots.


For a minute there were still good data. Files from:
RSA01-2025.04.02.04.13.44.629.tiq_spectrogram.npz
to:
RSA01-2025.04.02.04.14.47.805.tiq_spectrogram.npz



05:30 update from HKR: There is a problem with the ESR main power supply. We have to wait.
Attachment 1: 2025-04-02_05-18-32-212_tcl1146.png  293 kB  Uploaded Wed Apr 2 06:22:14 2025  | Hide | Hide all
2025-04-02_05-18-32-212_tcl1146.png
Attachment 2: 2025-04-02_05-21-50-973_tcl1146.png  356 kB  Uploaded Wed Apr 2 06:22:21 2025  | Hide | Hide all
2025-04-02_05-21-50-973_tcl1146.png
ELOG V3.1.5-fc6679b